OpenStudy (itsbribro):

What is the slope of the line that passes through the points (16, –1) and (–4, 10)? A. -20/11 B. -11/20 C. 11/20 D. 20/11

OpenStudy (anonymous):

I got -11/20

OpenStudy (marigirl):

gradient= (y2-y1)/(x2-x1) Gradient=(10--1)/(-4-16) M=11/-20 therefore its -11/20

OpenStudy (itsbribro):

that's what i thought it was

OpenStudy (itsbribro):

can you help me with one more please @OrangeMaster

OpenStudy (anonymous):

Sure

OpenStudy (itsbribro):

What is the slope of the line that passes through the points (14, 5) and (20, –4)? A. -3/2 B. -2/3 C. 2/3 D. 3/2

OpenStudy (anonymous):

Simplify 9/6

OpenStudy (itsbribro):

3/2 D

OpenStudy (marigirl):

its -9/6 which simplifies to -3/2

OpenStudy (itsbribro):

when i tried it by myself i got 9/6=3/2 like @OrangeMaster said

OpenStudy (marigirl):

What is the slope of the line that passes through the points (14, 5) and (20, –4) (x1,y1) and (x2,y2) m=(y2-y1)/(x2-x1) m=(-4-5)/(20-14) m=-9/6

OpenStudy (itsbribro):

ok

OpenStudy (itsbribro):

thank you @marigirl and @OrangeMaster for helping me!!

OpenStudy (marigirl):

all good :)